OpenStudy (anonymous):

find the indicated power using DeMoivres theorem (Express your fully simplified answer in the form a + bi.)(3+sqrt3i)^4

OpenStudy (anonymous):

you need two numbers, \(r=\sqrt{a^2+b^2}\) and \(\theta\) so write this in either exponential or trig form

OpenStudy (anonymous):

\(z=3+\sqrt{3}i\) and so \(a=3,b=\sqrt{3}, \sqrt{a^2+b^2}=\sqrt{3^2+\sqrt{3}^2}=\sqrt{12}\)

OpenStudy (anonymous):

as for \(\theta\) since you are in quadratant 1 you can use \(\tan^{-1}(\frac{b}{a})\) which in your case is \[\tan^{-1}(\frac{\sqrt{3}}{3})=\frac{\pi}{6}\]

OpenStudy (anonymous):

therefore you can write the number is trig form as \[\sqrt{12}\left(\cos(\frac{\pi}{6})+i\sin(\frac{\pi}{6})\right)\] unless you would prefer exponential from, not sure what class you are taking

OpenStudy (anonymous):

to raise the number to the 4th power, take the 4th power of \(\sqrt{12}\) which is \(144\) and then multiply the angle by 4 you good from there?

OpenStudy (anonymous):

Change (3+sqrt3i) into ploar Form

OpenStudy (anonymous):

*Polar

OpenStudy (anonymous):

\[r(\cos x + i sinx) ^n\] \[r^n (Cos nx + i Sin nx)\]

OpenStudy (anonymous):

how would I express this fully simplified in the form a + bi.)

OpenStudy (anonymous):

that is your job, but it is a relatively easy one when you multiply \(\frac{\pi}{6}\) by \(4\) you get \(\frac{2\pi}{3}\)

OpenStudy (anonymous):

so \[\left(\sqrt{12}\left(\cos(\frac{\pi}{6})+i\sin(\frac{\pi}{6})\right)\right)^4\] \[=144\left(\cos(\frac{2\pi}{3})+i\sin(\frac{2\pi}{3})\right)\] you only need to evaluate sine and cosine to finish

OpenStudy (anonymous):

144(-1/2+i (sqrt3)/2

OpenStudy (anonymous):

72+72isqrt3 is that right

OpenStudy (anonymous):

looks good except you dropped a minus sign

OpenStudy (anonymous):

-72+72isqrt3

OpenStudy (anonymous):

thank you so much satllite73